a pack contains n cards numbered 1 to n .Two consecutive cards are removed from the pack and sum of remaining card is 1224.If smaller of number on the removed card is k.Find value of k-20

Grade:11

1 Answers

Abhishek Singh
9 years ago
its iit 2013 integer type question
acc to given condition n
n(n+1)/2 -(k)-(k+1)=1224
get k= n(n+1)-2450 / 4
 since n should be such that n(n+1)>2450 and k is integer only n= 50 satisfies , 
then k= 25 and  k-20 = 5